the mass of the sphere is 3.2*10^-14 kg and it carries a net charge equal to that of 10 electrons.if the electronic charge is -1.6*10^-19 C and the acceleration of free fall is the field required to keep the sphere stationary is

Grade:11

1 Answers

VIBHANSHU GARG
6 years ago
Mass of sphere m=3.2×10^-14kgCharge on sphere q=10×1.6×10^-19=1.6×10^-18CAcceleration due to gravity g=10m/s^2As,qE=mg1.6×10^-18E=3.2×10^-14×10E=10^5N/CHence the electric field is 10^5N/C